# auxmulticols[1]

### v1

@@ 6.1. Recommendations for Regulators & CS Communities
Automated tools for assisting in data rights implementation and enforcement. The enforcement of laws poses significant technological challenges to companies and regulators (see C5, C7, and R3). Coupled with companies’ struggles to implement data rights and assess compliance status, these challenges underscore the need for automated compliance tools. Such tools could ease data rights implementation for companies and aid regulators in compliance monitoring and law enforcement. Consider, for instance, an automatic tool that evaluates a company’s compliance without manual intervention. This tool could help companies identify their compliance status and subsequently enhance data rights implementation. Also, some automatic tools can reduce the complexity of implementing a data right. Many researchers have pursued the development of such tools (235; 70; 12; 43; 136; 137; 194; 195; 217; 234; 230; 125; 127; 72), but efforts are disproportionally focused on the right to information. Despite significant progress, the real-world effectiveness of these tools remains questionable (see C4 and C5), suggesting that further work is needed to improve the data rights regime.
@@ 6.2. Privacy-Enhancing Technologies
The implementation of data rights presents numerous technical challenges. Furthermore, the unique nature of data amplifies these challenges, making it difficult for data owners to assert clear data rights. The current rights-based privacy approach often simplifies data as a property in the modern economy, which may not fully capture its characteristics. Privacy-Enhancing technologies can be instrumental in addressing these issues (73; 124).
@@
While these technologies have ideal properties to process sensitive data, they are still maturing and have not yet fulfilled their promise (206; 73; 124). In addition, the new data rights regime is not a panacea. Digital divide, psychological, and economic factors may still prevent users from fully exercising their data rights (see Section 3). Despite this, we posit that the adoption of these technologies could enhance the effectiveness of the rights-based approach.
@@ 7. Alternative Approaches to Data Rights
Throughout this paper, we have identified that one of the primary issues with rights-based approaches is the reliance on questionable user assumptions. Few users are willing to undertake the burden required to exercise their data rights. Consequently, an increasing number of scholars question the rights-based approach for this reason (201; 87; 115; 47; 26; 180; 227; 94; 78; 71). Many have proposed alternative solutions that supplement rights allocations, often by altering the relationship dynamics between data subjects (i.e., users) and controllers (i.e., companies).
@@
Some commentators favor the creation of an explicit duty of loyalty on data controllers towards data subjects, which is called “data fiduciary” or “trust” (87; 13; 14; 89; 65). This duty of loyalty would obligate controllers to design services and operate in ways that do not conflict with users’ interests. Duties represent potent obligations; if imposed within the information economy, they would challenge practices like “nudging,” wherein data is used to manipulate or unfairly target users. Loyalty obligations cannot be easily undone like rights-based approaches, as courts tend to critically construe waivers of duties against the company owing the obligation. Consequently, fine print, dark patterns, and manipulative interfaces would fail to undermine duties (87). However, duties do not offer a definitive solution for data power issues, as data controllers, unlike doctors and lawyers, must act in their economic self-interest. A duty of loyalty to data subjects is at odds with the longstanding consensus that companies’ primary moral imperative is to maximize shareholder value (115).
Rights-based approaches position individuals as pivotal actors in privacy power dynamics. Some commentators advocate for overturning this user-centricity, suggesting a collective-societal approach where “group privacy” becomes the central principle (180; 227; 94; 78; 71). Priscilla Regan was an early advocate for this collective viewpoint, foreseeing the power shifts resulting from data proliferation and arguing that collective privacy interests would better serve the political landscape (180).
@@ A.3. Regulators
Right to information (Arts. 13&14 of the GDPR)
Service providers that collect users’ data must provide data subjects with information about when and why they collect data, the way to exercise their data rights in the service, etc.
Right to access (Art. 15 of the GDPR)
The data subject should be able to access their personal data that companies process and ask for information on companies’ data processing.
Right to rectification (Art. 16 of the GDPR)
Subjects should be able to change and update their personal data that service providers store.
Right to erasure (Art. 17 of the GDPR)
Data subjects can erase their personal data by sending a request to the company under certain conditions (e.g., in the case where they withdraw their consent to data processing).
Right to restriction of processing (Art. 18 of the GDPR)
Users can ask companies to restrict the processing of their data under certain conditions (e.g., in the case where they contest the accuracy of the data).
Right to data portability (Art. 20 of the GDPR)
Users should be able to receive the personal data held by companies in a machine-readable form and send the data to other companies.
Right to object (Art. 21 of the GDPR)
Users can object to processing their personal data under certain conditions. The right to object to direct marketing is absolute.
Right to avoid automated decision-making (Art. 22 of the GDPR)
As the last data right of the GDPR, this right allows users to request not to be subject to a decision based solely on automatic processing, including profiling.
@@
Label
Description
Relevant ones of 14 factors
Policy-comp
Investigate whether privacy policies are written in compliance with GDPR
C4
Policy-read
Investigate how readable privacy policies are
C6
NewDesign
Propose a new system design to better support data rights
C6
Cookie-comp
Investigate whether cookie notices comply with GDPR
C4
Cookie-read
Investigate how readable cookie notices are
C6
User-know
User knowledge or awareness of data rights
U1
User-th
Users’ general thoughts on data rights (e.g., Perception, expectation, and willingness)
U2
User-behave
User behaviour and experience regarding data rights; in particular, how they handle their data in the real world
U3
User-effect
Data rights’ effects on users (e.g., their trust in companies, their privacy concerns, behaviour of sharing their personal information)
U4
Request-comp
Investigate whether service providers comply with GDPR by sending a data rights request to them
C4
System-comp
Investigate whether service providers comply with GDPR through a system analysis
C4
RWdata-comp
Investigate whether service providers comply with GDPR through an analysis of real-world data such as social media data and customers’ reviews
C4
Usab
Investigate usability of exercising data rights (except for an analysis of privacy policies)
C6
Dev-know
Developers’ (or companies’) knowledge or awareness of data rights
C1
Dev-th
Developers’ (or companies’) general thoughts on data rights (e.g., Perception, willingness, and attitude)
C2, C5
Dev-effort
Developers’ (or companies’) efforts and experience regarding the implementation of data rights;
C3, C4
Dev-effect
Data rights’ effects on companies (e.g., (dis)advantages of GDPR)
C8
Reg-effort
Regulator’ efforts to support data rights
R1, R2, R3
Law-th
Perception and thoughts of law experts
R1, R2
Infra
Investigate companies’ infrastructures that have changed to support data rights
C3
Dark
Dark patterns
C6
Flaw
Investigate whether there are flaws in the current implementation of data rights
C7
Fines
Investigate whether fines are sufficient to enforce the data rights
R3
@@
U1: User knowledge
(U1.2) The digital divide leads to a gap in GDPR data rights knowledge.
@@
U2: User perception
(U2.1) Many users express interest in having data rights across various contexts.
Users
(U2.2) However, many others don’t also feel the necessity of data rights due to high trust in companies that collect their data, indifference towards using their data, privacy fatigue caused, fear from learning about personal data processing, and a lack of understanding data rights.
@@
U3: User action
(U3.2) The evidence suggests that users don’t exercise their data rights very often in general.
@@
U4: Effect on users
(U4.1) Per the potential benefits reviewed, although having data rights can reduce users’ perceived privacy concerns in general, how data rights are implemented in practice has a significant influence as well. Moreover, the right to avoid automated decision-making little improves user perception.
@@
C1: Company knowledge
Many companies aren’t fully aware of GDPR data rights and don’t understand them well.
C2: Attitude towards compliance
(C2.1) Many developers remain unaware, indifferent, or skeptical about requirements to comply with the obligation to provide data rights.
@@
C3: Administrative efforts
Many companies have taken internal measures to facilitate their GDPR compliance, including employee training, changes in management, and procuring professional advice. However, some companies have decided to evade certain GDPR data rights.
C4: Compliance with the law
(C4.1) Many companies do not fully provide the right to information, and the level of compliance varies significantly across service applications.
@@
Companies or developers
C5: Significant difficulties
(C5.2) Implementing the rights to information and access is challenging due to complex data processing by third parties. Moreover, explaining ML algorithms to general users is difficult under the right to information, as these algorithms are often seen as black boxes and too technical for user comprehension. User identification is also considered a challenge in enforcing the right to access.
@@
C6: Usable design
(C6.2) But, there is conflicting evidence regarding the usability of data rights, as some companies successfully provide data rights with good usability. Therefore, the level of usability depends on the specific application or implementation.
@@
C7: New threats or flaws
The current implementation of the access right increases the risk of information leakage to unauthorized third parties.
C8: Effect on companies
(C8.1) The impact of data rights on compliant companies is unclear. While GDPR data rights can enhance internal data management, there is a possibility that users may hesitate to share their data under these rights, which hinders efficient data utilization by companies.
@@
R1: Monitoring
Regulators communicate with citizens, research, and hold conferences and seminars to understand the effects of data rights on users and companies. They are aware of some of the challenges in implementing data rights, but they also evaluate that the GDPR has overall empowered people.
Regulators
R2: Response and adjustments
Regulators provide additional guidelines to users and companies. However, it remains to be seen whether the proposed guidelines will be usable and achievable.
R3: Enforcement
The enforcement of GDPR data rights is currently lacking, with small fines, limited technical expertise, and funding challenges faced by regulatory authorities. Additionally, there is heterogeneity in enforcement across Europe.

### v2

@@ 6.1. Recommendations for Regulators & CS Communities
Automated tools for assisting in data rights implementation and enforcement. The enforcement of laws poses significant technological challenges to companies and regulators (see C5, C7, and R3). Coupled with companies’ struggles to implement data rights and assess compliance status, these challenges underscore the need for automated compliance tools. Such tools could ease data rights implementation for companies and aid regulators in compliance monitoring and law enforcement. Consider, for instance, an automatic tool that evaluates a company’s compliance without manual intervention. This tool could help companies identify their compliance status and subsequently enhance data rights implementation. Also, some automatic tools can reduce the complexity of implementing a data right. Many researchers have pursued the development of such tools (247; 75; 15; 46; 145; 146; 205; 206; 229; 246; 242; 133; 135; 77), but efforts are disproportionally focused on the right to information. Despite significant progress, the real-world effectiveness of these tools remains questionable (see C4 and C5), suggesting that further work is needed to improve the data rights regime.
@@ 6.2. Privacy-Enhancing Technologies
The implementation of data rights presents numerous technical challenges. Furthermore, the unique nature of data amplifies these challenges, making it difficult for data owners to assert clear data rights. The current rights-based privacy approach often simplifies data as a property in the modern economy, which may not fully capture its characteristics. Privacy-Enhancing technologies can be instrumental in addressing these issues (78; 132).
@@
While these technologies have ideal properties to process sensitive data, they are still maturing and have not yet fulfilled their promise (218; 78; 132). In addition, the new data rights regime is not a panacea. Digital divide, psychological, and economic factors may still prevent users from fully exercising their data rights (see Section 3). Despite this, we posit that the adoption of these technologies could enhance the effectiveness of the rights-based approach.
@@ 7. Alternative Approaches to Data Rights
Throughout this paper, we have identified that one of the primary issues with rights-based approaches is the reliance on questionable user assumptions. Few users are willing to undertake the burden required to exercise their data rights. Consequently, an increasing number of scholars question the rights-based approach for this reason (213; 92; 123; 50; 29; 191; 239; 99; 83; 76). Many have proposed alternative solutions that supplement rights allocations, often by altering the relationship dynamics between data subjects (i.e., users) and controllers (i.e., companies).
@@
Some commentators favor the creation of an explicit duty of loyalty on data controllers towards data subjects, which is called “data fiduciary” or “trust” (92; 16; 17; 94; 70). This duty of loyalty would obligate controllers to design services and operate in ways that do not conflict with users’ interests. Duties represent potent obligations; if imposed within the information economy, they would challenge practices like “nudging,” wherein data is used to manipulate or unfairly target users. Loyalty obligations cannot be easily undone like rights-based approaches, as courts tend to critically construe waivers of duties against the company owing the obligation. Consequently, fine print, dark patterns, and manipulative interfaces would fail to undermine duties (92). However, duties do not offer a definitive solution for data power issues, as data controllers, unlike doctors and lawyers, must act in their economic self-interest. A duty of loyalty to data subjects is at odds with the longstanding consensus that companies’ primary moral imperative is to maximize shareholder value (123).
Rights-based approaches position individuals as pivotal actors in privacy power dynamics. Some commentators advocate for overturning this user-centricity, suggesting a collective-societal approach where “group privacy” becomes the central principle (191; 239; 99; 83; 76). Priscilla Regan was an early advocate for this collective viewpoint, foreseeing the power shifts resulting from data proliferation and arguing that collective privacy interests would better serve the political landscape (191).
@@
8. Emerging Privacy Challenges in the Age of AI
The rapid proliferation of AI systems and the escalating race to train ever-larger models have surfaced many privacy concerns (153). Chief among them is the fact that training corpora—often scraped indiscriminately from the open web—routinely contain private or sensitive information. A growing body of work demonstrates that large models can memorize such data and later reveal it at inference time (253; 3; 113). Consequently, the composition of training datasets and the policies that govern their use matter profoundly.
Today, AI developers adopt heterogeneous privacy policies. OpenAI, for instance, retains user-model conversations for model improvement or training unless users explicitly opt out (168). Anthropic, by contrast, pledges not to incorporate user data into training by default (13). The absence of a harmonized industry standard and a comprehensive legal framework for training data governance exacerbates user data rights (54). As models continue to scale, the need for statutory clarity intensifies.
Legal safeguards, however, are merely one piece of the puzzle. If individuals wish to invoke the “right to erasure,” technical mechanisms must exist to remove their data from a trained model. Several approaches have been explored. Data-centric techniques attempt to delete sensitive text before training (118; 137), while others rely on differential privacy to provably limit memorization (67; 12; 255). Both strategies generally require retraining the model, incurring prohibitive computational costs at frontier scales. To avoid full retraining, researchers have proposed knowledge unlearning, which updates only a small subset of parameters to excise targeted knowledge (109). Although promising, current methods remain in their infancy and can trigger catastrophic forgetting or other safety failures (209). Further work—both technical and legal—is therefore essential.
Privacy risks are not limited to memorization. Even without data leakage, large multi-modal models can infer personally identifiable information from inputs (253). As capabilities grow, so too do privacy harms, underscoring the urgency of holistic solutions that span policy, model architecture, and deployment practices.
@@ A.3. Regulators
Right to information (Arts. 13&14 of the GDPR) Service providers that collect users’ data must provide data subjects with information about when and why they collect data, the way to exercise their data rights in the service, etc.
Right to access (Art. 15 of the GDPR) The data subject should be able to access their personal data that companies process and ask for information on companies’ data processing.
Right to rectification (Art. 16 of the GDPR) Subjects should be able to change and update their personal data that service providers store.
Right to erasure (Art. 17 of the GDPR) Data subjects can erase their personal data by sending a request to the company under certain conditions (e.g., in the case where they withdraw their consent to data processing).
Right to restriction of processing (Art. 18 of the GDPR) Users can ask companies to restrict the processing of their data under certain conditions (e.g., in the case where they contest the accuracy of the data).
Right to data portability (Art. 20 of the GDPR) Users should be able to receive the personal data held by companies in a machine-readable form and send the data to other companies.
Right to object (Art. 21 of the GDPR) Users can object to processing their personal data under certain conditions. The right to object to direct marketing is absolute.
Right to avoid automated decision-making (Art. 22 of the GDPR) As the last data right of the GDPR, this right allows users to request not to be subject to a decision based solely on automatic processing, including profiling.
@@
Label Description Relevant ones of 14 factors
Policy-comp Investigate whether privacy policies are written in compliance with GDPR C4
Policy-read Investigate how readable privacy policies are C6
NewDesign Propose a new system design to better support data rights C6
Cookie-comp Investigate whether cookie notices comply with GDPR C4
Cookie-read Investigate how readable cookie notices are C6
User-know User knowledge or awareness of data rights U1
User-th Users’ general thoughts on data rights (e.g., Perception, expectation, and willingness) U2
User-behave User behaviour and experience regarding data rights; in particular, how they handle their data in the real world U3
User-effect Data rights’ effects on users (e.g., their trust in companies, their privacy concerns, behaviour of sharing their personal information) U4
Request-comp Investigate whether service providers comply with GDPR by sending a data rights request to them C4
System-comp Investigate whether service providers comply with GDPR through a system analysis C4
RWdata-comp Investigate whether service providers comply with GDPR through an analysis of real-world data such as social media data and customers’ reviews C4
Usab Investigate usability of exercising data rights (except for an analysis of privacy policies) C6
Dev-know Developers’ (or companies’) knowledge or awareness of data rights C1
Dev-th Developers’ (or companies’) general thoughts on data rights (e.g., Perception, willingness, and attitude) C2, C5
Dev-effort Developers’ (or companies’) efforts and experience regarding the implementation of data rights; C3, C4
Dev-effect Data rights’ effects on companies (e.g., (dis)advantages of GDPR) C8
Reg-effort Regulator’ efforts to support data rights R1, R2, R3
Law-th Perception and thoughts of law experts R1, R2
Infra Investigate companies’ infrastructures that have changed to support data rights C3
Dark Dark patterns C6
Flaw Investigate whether there are flaws in the current implementation of data rights C7
Fines Investigate whether fines are sufficient to enforce the data rights R3
@@
U1: User knowledge (U1.2) The digital divide leads to a gap in GDPR data rights knowledge.
@@
U2: User perception (U2.1) Many users express interest in having data rights across various contexts.
Users (U2.2) However, many others don’t also feel the necessity of data rights due to high trust in companies that collect their data, indifference towards using their data, privacy fatigue caused, fear from learning about personal data processing, and a lack of understanding data rights.
@@
U3: User action (U3.2) The evidence suggests that users don’t exercise their data rights very often in general.
@@
U4: Effect on users (U4.1) Per the potential benefits reviewed, although having data rights can reduce users’ perceived privacy concerns in general, how data rights are implemented in practice has a significant influence as well. Moreover, the right to avoid automated decision-making little improves user perception.
@@
C1: Company knowledge Many companies aren’t fully aware of GDPR data rights and don’t understand them well.
C2: Attitude towards compliance (C2.1) Many developers remain unaware, indifferent, or skeptical about requirements to comply with the obligation to provide data rights.
@@
C3: Administrative efforts Many companies have taken internal measures to facilitate their GDPR compliance, including employee training, changes in management, and procuring professional advice. However, some companies have decided to evade certain GDPR data rights.
C4: Compliance with the law (C4.1) Many companies do not fully provide the right to information, and the level of compliance varies significantly across service applications.
@@
Companies or developers C5: Significant difficulties (C5.2) Implementing the rights to information and access is challenging due to complex data processing by third parties. Moreover, explaining ML algorithms to general users is difficult under the right to information, as these algorithms are often seen as black boxes and too technical for user comprehension. User identification is also considered a challenge in enforcing the right to access.
@@
C6: Usable design (C6.2) But, there is conflicting evidence regarding the usability of data rights, as some companies successfully provide data rights with good usability. Therefore, the level of usability depends on the specific application or implementation.
@@
C7: New threats or flaws The current implementation of the access right increases the risk of information leakage to unauthorized third parties.
C8: Effect on companies (C8.1) The impact of data rights on compliant companies is unclear. While GDPR data rights can enhance internal data management, there is a possibility that users may hesitate to share their data under these rights, which hinders efficient data utilization by companies.
@@
R1: Monitoring Regulators communicate with citizens, research, and hold conferences and seminars to understand the effects of data rights on users and companies. They are aware of some of the challenges in implementing data rights, but they also evaluate that the GDPR has overall empowered people.
Regulators R2: Response and adjustments Regulators provide additional guidelines to users and companies. However, it remains to be seen whether the proposed guidelines will be usable and achievable.
R3: Enforcement The enforcement of GDPR data rights is currently lacking, with small fines, limited technical expertise, and funding challenges faced by regulatory authorities. Additionally, there is heterogeneity in enforcement across Europe.